Carbex, Inc., produces cutlery sets out of high-quality wood and steel. The company makes a Standard set and a Deluxe set and sells them to retail department stores throughout the country. The Standard set sells for $72, and the Deluxe set sells for $87. The variable expenses associated with each set are given below.

Standard Deluxe
Variable production costs $ 21.00 $ 36.00
Sales commissions (21% of sales price) $ 15.12 $ 18.27

The company’s fixed expenses each month are:

Advertising $ 111,000
Depreciation $ 23,500
Administrative $ 66,000

Mary Parsons, the financial vice president, watches sales commissions carefully and has noted that they have risen steadily over the last year. For this reason, she was shocked to find that even though sales have increased, profits for the current month—May—are down substantially from April. Sales, in sets, for the last two months are given below:

Standard Deluxe Total
April 4,600 2,600 7,200
May 1,600 5,600 7,200

Required:

1-a. Prepare contribution format income statements for April.

1-b. Prepare contribution format income statements for May.

3-a. Compute the break-even point in dollar sales for April.

3-b. Would the break-even point in May be higher or lower than the break-even point in April?
Answer #1
1-a] Standard Deluxe Total
Sales price per unit $            72.00 $          87.00
Sales in units 4600 2600
Variable production costs per unit $            21.00 $          36.00
Sales commissions per unit [21%] $            15.12 $          18.27
Sales revenue $      3,31,200 $    2,26,200 $   5,57,400
Variable expenses:
Prodution $         96,600 $        93,600 $   1,90,200
Sales $         69,552 $        47,502 $   1,17,054
Total variable expenses $      1,66,152 $ 1,41,102 $   3,07,254
Contribution margin $      1,65,048 $        85,098 $   2,50,146
Fixed costs:
Advertising $   1,11,000
Depreciation $ 23,500
Administrative $ 66,000
Total fixed costs $   2,00,500
Net operating income $ 49,646
1-b] Standard Deluxe Total
Sales price per unit $            72.00 $          87.00
Sales in units 1600 5600
Variable production costs per unit $            21.00 $          36.00
Sales commissions per unit [21%] $            15.12 $          18.27
Sales revenue $      1,15,200 $    4,87,200 $   6,02,400
Variable expenses:
Prodution $         33,600 $    2,01,600 $   2,35,200
Sales $         24,192 $    1,02,312 $   1,26,504
Total variable expenses $         57,792 $    3,03,912 $   3,61,704
Contribution margin $         57,408 $    1,83,288 $   2,40,696
Fixed costs:
Advertising $   1,11,000
Depreciation $ 23,500
Administrative $ 66,000
Total fixed costs $   2,00,500
Net operating income $ 40,196
3-a] Sales mix for April 23 13
Contribution margin per unit $            35.88 $          32.73
Weighted average contribution margin ratio = 250146/557400 = 44.88%
BEP in $ for April = 200500/44.88% = $   4,46,774
3-b] Sales mix for May 2 7
Weighted average contribution margin ratio = 240696/602400 = 39.96%
The BEP in may will be higher as the weighted CM ratio is lower.
The BEP in $ would be 200500/39.96% = $   5,01,800